# Toward Conversational Hungarian Speech Recognition: Introducing the BEA-Large and BEA-Dialogue Datasets

### v1

@@ Abstract
The advancement of automatic speech recognition (ASR) has been largely enhanced by extensive datasets in high-resource languages, while languages such as Hungarian remain underrepresented due to limited spontaneous and conversational corpora. To address this gap, we introduce two new datasets – BEA-Large and BEA-Dialogue – constructed from the previously unprocessed portions of the Hungarian speech corpus named BEA. BEA-Large extends BEA-Base with 255 hours of spontaneous speech from 433 speakers, enriched with detailed segment-level metadata. BEA-Dialogue, comprising 85 hours of spontaneous conversations, is a Hungarian speech corpus featuring natural dialogues partitioned into speaker-independent subsets, supporting research in conversational ASR and speaker diarization. We establish reproducible baselines on these datasets using publicly available ASR models, with the fine-tuned Fast Conformer model achieving word error rates as low as 14.18% on spontaneous and 4.8% on repeated speech. Diarization experiments yield diarization error rates between 13.05% and 18.26%, providing reference points for future improvements. The results highlight the persistent difficulty of conversational ASR, particularly due to disfluencies, overlaps, and informal speech patterns. By releasing these datasets and baselines, we aim to advance Hungarian speech technology and offer a methodological framework for developing spontaneous and conversational benchmarks in other languages.
@@ 4.  Construction of BEA-Dialogue
# Segments 7,864 567 1,739
# Words 601,742 35,780 115,816
# Characters 3,763,514 226,291 727,901
Avg. # Speakers / Seg 1.79 1.93 1.62
Avg. Seg Duration [s] 30.72 26.79 28.70
Overlap Duration [h] 22.99 0.58 2.42
Total Duration [h] 67.11 4.22 13.86
@@ 6.  Baseline ASR results for BEA-Dialogue
whisper-large-v3 20.74 20.74 12.26 12.25 45.74 20.38 20.38 10.79 10.79 53.61
whisper-large-v2 19.60 19.60 9.86 9.85 46.67 22.31 22.31 11.35 11.35 53.86
whisper-medium 24.82 24.82 12.02 12.01 46.30 27.59 27.59 13.41 13.41 53.49
f_conformer_ctc 19.93 19.93 8.21 8.21 70.00 19.24 19.19 7.93 7.86 82.15
Table 6: ASR accuracy comparison on BEA-Dialogue.
@@ 8.  Conclusion
Our baseline experiments yield several key findings. Fine-tuned Fast Conformer models achieved WERs as low as 14.18% on spontaneous speech and 4.8% on repeated speech when trained on the combined datasets, highlighting the benefits of including data from both experiment leaders and discourse partners. In conversational ASR, the relatively high CER-to-WER ratio compared to monologue speech suggests that spontaneous dialogue introduces additional challenges—arising from phenomena such as overlapping speech, rapid turn-taking, and interaction-driven disfluency patterns that are not present in single-speaker settings. The speaker diarization baselines, with DERs ranging from 13.05% to 18.26%, establish solid reference points for future advancements.
@@ 9.  Bibliographical References
[…]

### v2

@@ Abstract
The advancement of automatic speech recognition (ASR) has been largely enhanced by extensive datasets in high-resource languages, while languages such as Hungarian remain underrepresented due to limited spontaneous and conversational corpora. To address this gap, we introduce two new datasets – BEA-Large and BEA-Dialogue – constructed from the previously unprocessed portions of the Hungarian speech corpus named BEA. BEA-Large extends BEA-Base with 255 hours of spontaneous speech from 433 speakers, enriched with detailed segment-level metadata. BEA-Dialogue, comprising 85 hours of spontaneous conversations, is a Hungarian speech corpus featuring natural dialogues partitioned into speaker-independent subsets, supporting research in conversational ASR and speaker diarization. We establish reproducible baselines on these datasets using publicly available ASR models, with the fine-tuned Fast Conformer model achieving word error rates as low as 14.18% on spontaneous and 4.8% on repeated speech. Diarization experiments yield diarization error rates between 12.46% and 17.40%, providing reference points for future improvements. The results highlight the persistent difficulty of conversational ASR, particularly due to disfluencies, overlaps, and informal speech patterns. By releasing these datasets and baselines, we aim to advance Hungarian speech technology and offer a methodological framework for developing spontaneous and conversational benchmarks in other languages.
@@ 4.  Construction of BEA-Dialogue
# Segments 9,179 577 1,906
# Words 532,732 34,056 105,472
# Characters 3,217,617 206,740 641,628
Avg. # Speakers / Seg 1.77 1.92 1.61
Avg. # Utterances / Seg 10.99 8.68 9.74
Avg. Seg Duration [s] 26.23 26.31 26.09
Overlap Duration [h] 3.28 0.29 0.40
Total Duration [h] 66.87 4.22 13.81
@@ 6.  Baseline ASR results for BEA-Dialogue
whisper-large-v3 21.19 21.04 12.74 12.56 22.21 22.13 12.27 12.18
whisper-large-v2 19.65 19.42 9.84 9.58 24.50 24.42 13.13 13.05
whisper-medium 25.45 25.27 12.61 12.42 29.21 29.12 14.71 14.63
f_conformer_ctc 19.69 19.53 7.95 7.78 69.32 20.56 20.44 9.11 9.00 82.16
Table 6: ASR accuracy (%) comparison on BEA-Dialogue.
@@ 8.  Conclusion
Our baseline experiments yield several key findings. Fine-tuned Fast Conformer models achieved WERs as low as 14.18% on spontaneous speech and 4.8% on repeated speech when trained on the combined datasets, highlighting the benefits of including data from both experiment leaders and discourse partners. In conversational ASR, the relatively high CER-to-WER ratio compared to monologue speech suggests that spontaneous dialogue introduces additional challenges—arising from phenomena such as overlapping speech, rapid turn-taking, and interaction-driven disfluency patterns that are not present in single-speaker settings. The speaker diarization baselines, with DERs ranging from 12.46% to 17.40%, establish solid reference points for future advancements.
@@ 9.  Bibliographical References
  • L. Besacier, E. Barnard, A. Karpov, and T. Schultz (2014) Automatic speech recognition for under-resourced languages: a survey. Speech Communication 56, pp. 85–100. External Links: ISSN 0167-6393, Document, Link Cited by: §1.
  • H. Bredin (2023) Pyannote.audio 2.1 speaker diarization pipeline: principle, benchmark, and recipe. In Interspeech 2023, pp. 1983–1987. External Links: Document, ISSN 2958-1796 Cited by: §7.
  • N. Kanda, Y. Gaur, X. Wang, Z. Meng, and T. Yoshioka (2020) Serialized output training for end-to-end overlapped speech recognition. In Interspeech 2020, pp. 2797–2801. External Links: Document, ISSN 2958-1796 Cited by: §6.
  • O. Kuchaiev, J. Li, H. Nguyen, O. Hrinchuk, R. Leary, B. Ginsburg, S. Kriman, S. Beliaev, V. Lavrukhin, J. Cook, P. Castonguay, M. Popova, J. Huang, and J. M. Cohen (2019) NeMo: a toolkit for building ai applications using neural modules. External Links: 1909.09577, Link Cited by: §5.
  • J. Linke, P. N. Garner, G. Kubin, and B. Schuppler (2022) Conversational speech recognition needs data? experiments with Austrian German. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, N. Calzolari, F. Béchet, P. Blache, K. Choukri, C. Cieri, T. Declerck, S. Goggi, H. Isahara, B. Maegaard, J. Mariani, H. Mazo, J. Odijk, and S. Piperidis (Eds.), Marseille, France, pp. 4684–4691. External Links: Link Cited by: §2.
  • P. Mihajlik, K. Mády, A. Kohári, F. S. Fruzsina, G. Kiss, T. E. Gráczi, and A. S. Doğruöz (2024) Is spoken Hungarian low-resource?: a quantitative survey of Hungarian speech data sets. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), N. Calzolari, M. Kan, V. Hoste, A. Lenci, S. Sakti, and N. Xue (Eds.), Torino, Italia, pp. 9382–9388. External Links: Link Cited by: §1, §2.
  • P. Mihajlik, Z. Tuske, B. Tarján, B. Németh, and T. Fegyó (2010) Improved recognition of spontaneous hungarian speech—morphological and acoustic modeling techniques for a less resourced task. IEEE Transactions on Audio, Speech, and Language Processing 18 (6), pp. 1588–1600. External Links: Document Cited by: §1.
  • T. J. Park, I. Medennikov, K. Dhawan, W. Wang, H. Huang, N. R. Koluguri, K. C. Puvvada, J. Balam, and B. Ginsburg (2024) Sortformer: a novel approach for permutation-resolved speaker supervision in speech-to-text systems. External Links: Link Cited by: §7.
  • A. Radford, J. W. Kim, T. Xu, G. Brockman, C. McLeavey, and I. Sutskever (2022) Robust speech recognition via large-scale weak supervision. External Links: 2212.04356, Link Cited by: §5.
  • M. Ravanelli, T. Parcollet, A. Moumen, S. de Langen, C. Subakan, P. Plantinga, Y. Wang, P. Mousavi, L. D. Libera, A. Ploujnikov, F. Paissan, D. Borra, S. Zaiem, Z. Zhao, S. Zhang, G. Karakasidis, S. Yeh, P. Champion, A. Rouhe, R. Braun, F. Mai, J. Zuluaga-Gomez, S. M. Mousavi, A. Nautsch, H. Nguyen, X. Liu, S. Sagar, J. Duret, S. Mdhaffar, G. Laperrière, M. Rouvier, R. D. Mori, and Y. Estève (2024) Open-source conversational ai with speechbrain 1.0. Journal of Machine Learning Research 25 (333). External Links: Link Cited by: §5.
  • M. Ravanelli, T. Parcollet, P. Plantinga, A. Rouhe, S. Cornell, L. Lugosch, C. Subakan, N. Dawalatabad, A. Heba, J. Zhong, J. Chou, S. Yeh, S. Fu, C. Liao, E. Rastorgueva, F. Grondin, W. Aris, H. Na, Y. Gao, R. D. Mori, and Y. Bengio (2021) SpeechBrain: a general-purpose speech toolkit. Note: arXiv:2106.04624 External Links: 2106.04624 Cited by: §5.